QUESTION 28 Let f(x) = 2 - X and g(x)= 3- X, find (fog)(2) O undefined 02 Orz O,
Answer #1

Let f(x) = 2 – X and g(x)=/3 - X, find (fog)(2) O undefined O2 ora 0, (fog) (2)=f(g(2)) 9123 = √3-2 g12) = 1 so, flge)) = f(1